2 buses are approaching each other on a long straight road, at a constant speed of 50 km/hr each. When the buses are 10 km apart, a fly in front of one bus flies at a speed of 100 km/hr towards the other bus. When it arrives just in front of that bus, it turns back towards the first bus. This goes on, till it is finally sandwiched between the buses. Find the total distance travelled by the fly
 

Grade:10

3 Answers

Aditya Gupta
4 years ago
this is an ez ques. 
the time taken for the buses to meet= 10/100= 0.1hr
now, the total distance travelled by the fly= time for which it will travel* speed at which it is travelling
= 0.1*100
10km
kindly approve :)
Vinay
4 years ago
Distance=speed *time
Time =10/100=0.1 hour
The mosquito total time travelled is equal to the time where the 2 buses cross each other. 
D=100*0.1=10 km
Khimraj
4 years ago
Time =10/100=0.1 hour
distance = 100*0.1 = 10 km
.................................................................................................................................